OpenStudy (anonymous):

simplify the expression then add its the very first part I am not getting how to simplify it \[\sqrt[3]{54m^4}+m \sqrt[3]{16m} \]

OpenStudy (anonymous):

Hi satellite73

OpenStudy (anonymous):

first of all \[54=27\times 2\] so \[\sqrt[3]{54}=\sqrt[3]{27\times 2}=\sqrt[3]{27}\times \sqrt[3]{2}=3\sqrt[3]{2}\]

OpenStudy (anonymous):

hi so we can start to rewrite the first term

OpenStudy (anonymous):

also \[\sqrt[3]{m^4}=\sqrt[3]{m^3\times m}=\sqrt[3]{m^3}\times \sqrt[3]{m}=m\sqrt[3]{m}\]

OpenStudy (anonymous):

putting these together we get \[\sqrt[3]{54m^4}=3m\sqrt[3]{2m}\]

OpenStudy (anonymous):

for the second term, \(16=8\times 2\) so \[\sqrt[3]{16}=\sqrt[3]{8}\times \sqrt[3]{2}=2\sqrt[3]{2}\]

OpenStudy (anonymous):

so first term in simplest radical form is \[3m\sqrt[3]{2m}\] and second term in simplest radical form is \[2m\sqrt[3]{2m}\] and since these are like terms you can combine them as follows: \[3m\sqrt[3]{2m}+2m\sqrt[3]{2m}=5m\sqrt[3]{2m}\]

OpenStudy (anonymous):

you got this?

OpenStudy (anonymous):

ok I think I finally got it, I was making sure before I typed anything, you helped me break it down so that I understood how to do it. OMG TY sooooo much

OpenStudy (anonymous):

yw glad to help